OpenStudy (anonymous):

Need some help with Variation of Parameters please: y''+2y'+2y=e^(-x) cscx

OpenStudy (anonymous):

@SithsAndGiggles the problem shows as: \[y''+2y'+2y=e^{-x}cscx\]

OpenStudy (anonymous):

Characteristic solutions: \[r^2+2r+2=0\\ r=\frac{-2\pm2i}{2}=-1\pm i\] yielding \[y_c=C_1e^{-x}\cos x+C_2e^{-x}\sin x\] If I'm not mistaken... This makes \[y_1=e^{-x}\cos x\\ y_2=e^{-x}\sin x\]

OpenStudy (anonymous):

I am a bit confused on what Yc would be.. \[Yc=e^x(c _{1} \cos x+c _{2} sinx) \]?

OpenStudy (anonymous):

My bad, you're right

OpenStudy (anonymous):

Im working on the Wronskian..

OpenStudy (anonymous):

Okay, so now you can use the formulas for VoP: \[u_1=-\int\frac{e^{-x}\sin xe^{-x}\csc x}{W(y_1,y_2)}~dx=-\int\frac{e^{-2x}}{W(y_1,y_2)}~dx\\ u_2=\int\frac{e^{-x}\cos xe^{-x}\csc x}{W(y_1,y_2)}~dx=\int\frac{e^{-2x}\cot x}{W(y_1,y_2)}~dx\] where \[\begin{align*}W(y_1,y_2)&=\begin{vmatrix}e^{-x}\cos x&e^{-x}\sin x\\ -e^{-x}\cos x-e^{-x}\sin x&-e^{-x}\sin x+e^{-x}\cos x\end{vmatrix}\\\\ &=e^{-2x}\bigg[-\cos x\sin x+\cos^2 x+\sin x\cos x+\sin^2 x\bigg]\\ &=e^{-2x} \end{align*}\] Hmm, convenient.

OpenStudy (anonymous):

Would the Wronskian not equal \[2e^{-2x}\]?

OpenStudy (anonymous):

Ok i see, my bad you're right..

OpenStudy (anonymous):

Awesome! Thank you so much!

OpenStudy (anonymous):

You're welcome!